# Optimal frames for Phase Retrieval from Edge Vectors of Optimal Polygons

### v1

@@ Abstract.
This paper aims to characterize the optimal frame for phase retrieval, defined as the frame whose condition number for phase retrieval attains its minimal value. In the context of the two-dimensional real case, we reveal the connection between optimal frames for phase retrieval and the perimeter-maximizing isodiametric problem, originally proposed by Reinhardt in 1922. Our work establishes that every optimal solution to the perimeter-maximizing isodiametric problem inherently leads to an optimal frame in ${\mathbb{R}}^{2}$. By recasting the optimal polygons problem as one concerning the discrepancy of roots of unity, we characterize all optimal polygons. Building upon this connection, we then characterize all optimal frames with $m$ vectors in ${\mathbb{R}}^{2}$ for phase retrieval when $m\geq 3$ has an odd factor. As a key corollary, we show that the harmonic frame $E_{m}$ is not optimal for any even integer $m\geq 4$. This finding disproves a conjecture proposed by Xia, Xu, and Xu (Math. Comp., 90(356): 2931-2960). Previous work has established that the harmonic frame $E_{m}\subset{\mathbb{R}}^{2}$ is indeed optimal when $m$ is an odd integer.
@@
Zhiqiang Xu is supported by the National Science Fund for Distinguished Young Scholars (12025108), NSFC grant (12471361, 12021001, 12288201) and National Key R&D Program of China (2023YFA1009401).
@@ 1.1. Stability of phase retrieval
Let $\bm{A}=[\bm{a}_{1},\ldots,\bm{a}_{m}]^{T}\in\mathbb{H}^{m\times d}$ be a given measurement matrix, where each $\bm{a}_{j}\in\mathbb{H}^{d}$ and $\mathbb{H}\in\{\mathbb{R},\mathbb{C}\}$. The phase retrieval problem seeks to recover an unknown signal $\bm{x}\in\mathbb{H}^{d}$ from its phaseless measurements $\left|\left\langle\bm{a}_{j},\bm{x}\right\rangle\right|$, $j=1,\ldots,m$. Define the nonlinear measurement map $\Phi_{\bm{A}}:\mathbb{H}^{d}\rightarrow\mathbb{R}_{+}^{m}$ by
$\Phi_{\bm{A}}(\bm{x})=|\bm{A}\bm{x}|:=\left(\left|\left\langle\bm{a}_{1},\bm{x}\right\rangle\right|,\left|\left\langle\bm{a}_{2},\bm{x}\right\rangle\right|,\ldots,\left|\left\langle\bm{a}_{m},\bm{x}\right\rangle\right|\right)^{T}\in\mathbb{R}_{+}^{m}$
@@
A matrix $\bm{A}$ is said to have the phase retrieval property if $|\bm{A}\bm{x}|=|\bm{A}\bm{y}|$ implies $\bm{x}=c\cdot\bm{y}$ for some $c\in\mathbb{H}$ with $|c|=1$. It is known that $m\geq 2d-1$ (when $\mathbb{H}=\mathbb{R}$) or $m\geq 4d-4$ (when $\mathbb{H}=\mathbb{C}$) generic measurements suffice for phase retrieval property [5, 8, 13, 31].
@@
A key metric for evaluating the robustness of the phase retrieval is the condition number. For any $\bm{x},\bm{y}\in\mathbb{H}^{d}$, we define the distance between $\bm{x}$ and $\bm{y}$ as
$\operatorname{dist}_{\mathbb{H}}(\bm{x},\bm{y}):=\min\left\{\|\bm{x}-c\cdot\bm{y}\|_{2}:c\in\mathbb{H},|c|=1\right\},$
@@
where $\|\cdot\|_{2}$ represents the Euclidean norm. The condition number $\beta_{\bm{A}}^{\mathbb{H}}$ of a given measurement matrix ${\bm{A}}\in\mathbb{H}^{m\times d}$ is defined as
$\beta_{\bm{A}}^{\mathbb{H}}:=\frac{U_{\bm{A}}^{\mathbb{H}}}{L_{\bm{A}}^{\mathbb{H}}},$
@@
where $L_{\bm{A}}^{\mathbb{H}}$ and $U_{\bm{A}}^{\mathbb{H}}$ represent the lower and upper Lipschitz constants of the map $\Phi_{\bm{A}}$, i.e.,
@@
$L_{\bm{A}}^{\mathbb{H}}:=\inf_{\begin{subarray}{c}\bm{x},\bm{y}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\bm{x},\bm{y})\neq 0\end{subarray}}\frac{\||\bm{A}\bm{x}|-|\bm{A}\bm{y}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\bm{x},\bm{y})}\quad\text{ and }\quad U_{\bm{A}}^{\mathbb{H}}:=\sup_{\begin{subarray}{c}\bm{x},\bm{y}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\bm{x},\bm{y})\neq 0\end{subarray}}\frac{\||\bm{A}\bm{x}|-|\bm{A}\bm{y}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\bm{x},\bm{y})}.$ (1.1)
@@
A smaller value of $\beta_{\bm{A}}^{\mathbb{H}}$ entails that $\Phi_{\bm{A}}$ behaves more like a near-isometry, indicating greater stability against the measurement noise. If $\bm{A}$ lacks the phase retrieval property, then $L_{\bm{A}}^{\mathbb{H}}=0$ and $\beta_{\bm{A}}^{\mathbb{H}}=+\infty$. We say that a matrix $\bm{A}\in\mathbb{H}^{m\times d}$ has the minimal condition number if $\beta_{\bm{A}}^{\mathbb{H}}=\min_{\bm{M}\in\mathbb{H}^{m\times d}}\beta_{\bm{M}}^{\mathbb{H}}$. To simplify the notation, we often omit the superscript in $\beta_{\bm{A}}^{\mathbb{H}}$, determining whether $\beta_{\bm{A}}^{\mathbb{H}}$ is defined over the real or complex field based on whether the matrix $\bm{A}$ is real or complex. In the same way, we drop the superscripts in $L_{\bm{A}}^{\mathbb{H}},U_{\bm{A}}^{\mathbb{H}}$, and we omit the subscript in $\operatorname{dist}_{\mathbb{H}}(\bm{x},\bm{y})$ when the underlying field is clear from context.
@@
Recently, the authors in [32] derive the first universal constant lower bound on the condition number $\beta_{\bm{A}}^{\mathbb{H}}$ for all $\bm{A}\in\mathbb{H}^{m\times d}$, namely,
@@
$\beta_{\bm{A}}^{\mathbb{H}}\geq\beta_{0}^{\mathbb{H}}:=\begin{cases}\sqrt{\frac{\pi}{\pi-2}}\approx 1.659&\text{ if }\mathbb{H}=\mathbb{R},\\ \sqrt{\frac{4}{4-\pi}}\approx 2.159&\text{ if }\mathbb{H}=\mathbb{C}.\end{cases}$ (1.2)
@@
The condition number of a standard Gaussian matrix in $\mathbb{H}^{m\times d}$ asymptotically matches the lower bound $\beta_{0}^{\mathbb{H}}$ as $m\rightarrow\infty$ [32]. Furthermore, in the real case $\mathbb{H}=\mathbb{R}$, they show that for all $\bm{A}\in\mathbb{R}^{m\times d}$,
$\beta_{\bm{A}}^{\mathbb{R}}\geq\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$ (1.3)
@@
$\beta_{\bm{E}_{m}}^{\mathbb{R}}=\begin{cases}\frac{1}{\sqrt{1-\frac{2}{m\cdot\sin\frac{\pi}{m}}}}&\text{if $m\geq 3$ is even,}\\ \frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}&\text{if $m\geq 3$ is odd,}\end{cases}$ (1.4)
where $\bm{E}_{m}$ is the harmonic frame in ${\mathbb{R}}^{2}$ which is defined by
$\bm{E}_{m}:=\left(\begin{array}[]{cccc}1&\cos\frac{1}{m}\pi&\cdots&\cos\frac{m-1}{m}\pi\\ 0&\sin\frac{1}{m}\pi&\cdots&\sin\frac{m-1}{m}\pi\end{array}\right)^{T}\in\mathbb{R}^{m\times 2}.$ (1.5)
Combining (1.3) and (1.4), we have $\beta_{\bm{E}_{m}}^{\mathbb{R}}=\min_{\bm{A}\in\mathbb{R}^{m\times 2}}\beta_{\bm{A}}^{\mathbb{R}}$ for each odd integer $m\geq 3$. The following conjecture is proposed in [32] for each even integer $m\geq 4$.
@@
$\beta_{\bm{E}_{m}}^{\mathbb{R}}=\min_{\bm{A}\in\mathbb{R}^{m\times 2}}\beta_{\bm{A}}^{\mathbb{R}}.$
In this paper, we will show that minimizing the condition number of $\bm{A}\in\mathbb{R}^{m\times 2}$ can be transformed into solving the perimeter-maximizing isodiametric problem proposed by Reinhardt in 1922 [25]. We prove that each optimal solution to the perimeter-maximizing isodiametric problem gives rise to a matrix with the minimal condition number. Based on that, we can show that Conjecture 1.1 is false. For instance, when $m=4$, we will demonstrate in Remark 1.3 that
$\beta_{\bm{E}_{4}}^{\mathbb{R}}={\sqrt{2+\sqrt{2}}}\approx 1.84>\beta_{\bm{E}_{4}^{\prime}}^{\mathbb{R}}=\min_{\bm{M}\in\mathbb{R}^{4\times 2}}\beta_{\bm{M}}^{\mathbb{R}}=\sqrt{1+\frac{\sqrt{2}}{2}+\frac{\sqrt{6}}{2}}\approx 1.71,$
@@
$\bm{E}_{4}^{\prime}=\left(\begin{array}[]{cccc}-1&\cos\frac{2\pi}{3}&\sqrt{2\sin\frac{\pi}{12}}\cos\frac{3\pi}{8}&\sqrt{2\sin\frac{\pi}{12}}\cos\frac{7\pi}{24}\\ 0&\sin\frac{2\pi}{3}&\sqrt{2\sin\frac{\pi}{12}}\sin\frac{3\pi}{8}&\sqrt{2\sin\frac{\pi}{12}}\sin\frac{7\pi}{24}\end{array}\right)^{T}.$ (1.6)
@@ 1.2. The perimeter-maximizing isodiametric problem for convex polygons
The perimeter-maximizing isodiametric problem has a long and rich history. It was first studied by Reinhardt in 1922 [25], and has been furthered investigated by Vincze [30] in 1950 and Datta [14] in 1997 (see also [15, 16, 24, 9]). The convex polygon in $\mathbb{R}^{2}$ with $m$ edges is called the convex $m$-gon. The diameter of a polygon $P$, denoted by $\operatorname{diam}({P})$, is the largest distance between any pair of its vertices, i.e., $\operatorname{diam}({P}):=\max_{\bm{x},\bm{y}\,\in{P}}{\|\bm{x}-\bm{y}\|_{2}}$. The perimeter of a polygon $P$, denoted by $\operatorname{perim}({P})$, is the sum of the length of its edges. The perimeter-maximizing isodiametric problem asks:
@@
A finite set in $\mathbb{R}^{d}$ is called a frame if the vectors in this set can span $\mathbb{R}^{d}$. We say that a frame $\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}\subset\mathbb{R}^{d}$ is optimal if its associated matrix $\bm{A}=[\bm{a}_{1},\ldots,\bm{a}_{m}]^{T}\in\mathbb{R}^{m\times d}$ has the minimal condition number, i.e., $\beta_{\bm{A}}^{\mathbb{R}}=\min_{\bm{M}\in\mathbb{R}^{m\times d}}\beta_{\bm{M}}^{\mathbb{R}}$. A frame $\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}\subset\mathbb{R}^{d}$ is called a tight frame if $\sum_{i=1}^{m}\bm{a}_{i}\bm{a}_{i}^{T}=C\cdot\bm{I}_{d}$ for some constant $C>0$.
@@
$f(\bm{a})\,\,=\,\,t^{2}\cdot(\cos 2\phi,\sin 2\phi)^{T}$ (1.10)
for any $\bm{a}=t\cdot(\cos\phi,\sin\phi)^{T}\in\mathcal{T}$.
@@
We will now show that the map $f:\mathcal{T}\to\mathbb{R}^{2}$ is a bijection. The bijectivity can be established by considering two cases. Firstly, for the zero vector, it holds that $\bm{a}=\bm{0}$ if and only if $f(\bm{a})=\bm{0}$. Secondly, for any nonzero vector $\bm{e}_{0}\in\mathbb{R}^{2}$, it possesses a unique representation in the form $\bm{e}_{0}=t_{0}\cdot(\cos{2\phi_{0}},\sin{2\phi_{0}})^{T}$, where $t_{0}>0$ and $\phi_{0}\in(0,\pi]$. Given this, the unique pre-image $\bm{a}_{0}\in\mathcal{T}$ such that $f(\bm{a}_{0})=\bm{e}_{0}$ is given by $\bm{a}_{0}=\sqrt{t_{0}}\cdot(\cos{\phi_{0}},\sin{\phi_{0}})^{T}$. Therefore, $f$ is both injective and surjective, implying it is a bijection. This ensures that the inverse map $f^{-1}:\mathbb{R}^{2}\to\mathcal{T}$ is well-defined, and is explicitly given by $f^{-1}(\bm{e}_{0})=\bm{a}_{0}$.

### v2

@@ Abstract.
This paper aims to characterize the optimal frame for phase retrieval, defined as the frame whose condition number for phase retrieval attains its minimal value. In the context of the two-dimensional real case, we reveal the connection between optimal frames for phase retrieval and the perimeter-maximizing isodiametric problem, originally proposed by Reinhardt in 1922. Our work establishes that every optimal solution to the perimeter-maximizing isodiametric problem inherently leads to an optimal frame in ${\mathbb{R}}^{2}$. By recasting the optimal polygons problem as one concerning the discrepancy of roots of unity, we characterize all optimal polygons. Building upon this connection, we then characterize all optimal frames with $m$ vectors in ${\mathbb{R}}^{2}$ for phase retrieval when $m\geq 3$ has an odd factor. As a key corollary, we show that the harmonic frame $E_{m}$ is not optimal for any even integer $m\geq 4$. This finding disproves a conjecture proposed by Xia, Xu, and Xu (Math. Comp., 94(356): 2931-2960). Previous work has established that the harmonic frame $E_{m}\subset{\mathbb{R}}^{2}$ is indeed optimal when $m$ is an odd integer.
@@
Key words and phrases:
Phase retrieval, stability, polygons, frames.
2020 Mathematics Subject Classification:
Primary 94A15, 46C05, 52B60; Secondary 94A12, 49N45, 52A10
The first author was supported by the National Science Fund for Distinguished Young Scholars (12025108) and NSFC grant (12471361, 12021001, 12288201).
The third author is the corresponding author.
@@ 1.1. Stability of phase retrieval
Let $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}\in\mathbb{H}^{m\times d}$ be a given measurement matrix, where each $\boldsymbol{a}_{j}\in\mathbb{H}^{d}$ and $\mathbb{H}\in\{\mathbb{R},\mathbb{C}\}$. The phase retrieval problem seeks to recover an unknown signal $\boldsymbol{x}\in\mathbb{H}^{d}$ from its phaseless measurements $\left|\left\langle\boldsymbol{a}_{j},\boldsymbol{x}\right\rangle\right|$, $j=1,\ldots,m$. Define the nonlinear measurement map $\Phi_{\boldsymbol{A}}:\mathbb{H}^{d}\rightarrow\mathbb{R}_{+}^{m}$ by
$\Phi_{\boldsymbol{A}}(\boldsymbol{x})=|\boldsymbol{A}\boldsymbol{x}|:=\left(\left|\left\langle\boldsymbol{a}_{1},\boldsymbol{x}\right\rangle\right|,\left|\left\langle\boldsymbol{a}_{2},\boldsymbol{x}\right\rangle\right|,\ldots,\left|\left\langle\boldsymbol{a}_{m},\boldsymbol{x}\right\rangle\right|\right)^{T}\in\mathbb{R}_{+}^{m}$ (1.1)
@@
A matrix $\boldsymbol{A}$ is said to have the phase retrieval property if $|\boldsymbol{A}\boldsymbol{x}|=|\boldsymbol{A}\boldsymbol{y}|$ implies $\boldsymbol{x}=c\cdot\boldsymbol{y}$ for some $c\in\mathbb{H}$ with $|c|=1$. It is known that $m\geq 2d-1$ (when $\mathbb{H}=\mathbb{R}$) or $m\geq 4d-4$ (when $\mathbb{H}=\mathbb{C}$) generic measurements suffice for phase retrieval property [5, 8, 13, 31].
@@
A key metric for evaluating the robustness of the phase retrieval is the condition number. For any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d}$, we define the distance between $\boldsymbol{x}$ and $\boldsymbol{y}$ as
$\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y}):=\min\left\{\|\boldsymbol{x}-c\cdot\boldsymbol{y}\|_{2}:c\in\mathbb{H},|c|=1\right\},$
@@
where $\|\cdot\|_{2}$ represents the Euclidean norm. The condition number $\beta_{\boldsymbol{A}}^{\mathbb{H}}$ of a given measurement matrix ${\boldsymbol{A}}\in\mathbb{H}^{m\times d}$ is defined as
$\beta_{\boldsymbol{A}}^{\mathbb{H}}:=\frac{U_{\boldsymbol{A}}^{\mathbb{H}}}{L_{\boldsymbol{A}}^{\mathbb{H}}},$
@@
where $L_{\boldsymbol{A}}^{\mathbb{H}}$ and $U_{\boldsymbol{A}}^{\mathbb{H}}$ represent the optimal lower and upper Lipschitz constants of the map $\Phi_{\boldsymbol{A}}$, i.e.,
@@
$L_{\boldsymbol{A}}^{\mathbb{H}}:=\inf_{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})\neq 0\end{subarray}}\frac{\||\boldsymbol{A}\boldsymbol{x}|-|\boldsymbol{A}\boldsymbol{y}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})}\quad\text{ and }\quad U_{\boldsymbol{A}}^{\mathbb{H}}:=\sup_{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})\neq 0\end{subarray}}\frac{\||\boldsymbol{A}\boldsymbol{x}|-|\boldsymbol{A}\boldsymbol{y}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})}.$
@@
A smaller value of $\beta_{\boldsymbol{A}}^{\mathbb{H}}$ entails that $\Phi_{\boldsymbol{A}}$ behaves more like a near-isometry, indicating greater stability against the measurement noise. If $\boldsymbol{A}$ lacks the phase retrieval property, then $L_{\boldsymbol{A}}^{\mathbb{H}}=0$ and $\beta_{\boldsymbol{A}}^{\mathbb{H}}=+\infty$. We say that a matrix $\boldsymbol{A}\in\mathbb{H}^{m\times d}$ has the minimal condition number if $\beta_{\boldsymbol{A}}^{\mathbb{H}}=\min_{\boldsymbol{M}\in\mathbb{H}^{m\times d}}\beta_{\boldsymbol{M}}^{\mathbb{H}}$. To simplify the notation, we often omit the superscript in $\beta_{\boldsymbol{A}}^{\mathbb{H}}$, determining whether $\beta_{\boldsymbol{A}}^{\mathbb{H}}$ is defined over the real or complex field based on whether the matrix $\boldsymbol{A}$ is real or complex. In the same way, we drop the superscripts in $L_{\boldsymbol{A}}^{\mathbb{H}},U_{\boldsymbol{A}}^{\mathbb{H}}$, and we omit the subscript in $\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})$ when the underlying field is clear from context.
@@
Recently, the authors in [32] derive the first universal constant lower bound on the condition number $\beta_{\boldsymbol{A}}^{\mathbb{H}}$ for all $\boldsymbol{A}\in\mathbb{H}^{m\times d}$, namely,
@@
$\beta_{\boldsymbol{A}}^{\mathbb{H}}\geq\beta_{0}^{\mathbb{H}}:=\begin{cases}\sqrt{\frac{\pi}{\pi-2}}\approx 1.659&\text{ if }\mathbb{H}=\mathbb{R},\\ \sqrt{\frac{4}{4-\pi}}\approx 2.159&\text{ if }\mathbb{H}=\mathbb{C}.\end{cases}$ (1.2)
@@
The condition number of a standard Gaussian matrix in $\mathbb{H}^{m\times d}$ asymptotically matches the lower bound $\beta_{0}^{\mathbb{H}}$ as $m\rightarrow\infty$ [32]. Furthermore, in the real case $\mathbb{H}=\mathbb{R}$, they show that for all $\boldsymbol{A}\in\mathbb{R}^{m\times d}$,
$\beta_{\boldsymbol{A}}^{\mathbb{R}}\geq\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$ (1.3)
@@
$\beta_{\boldsymbol{E}_{m}}^{\mathbb{R}}=\begin{cases}\frac{1}{\sqrt{1-\frac{2}{m\cdot\sin\frac{\pi}{m}}}}&\text{if $m\geq 3$ is even,}\\ \frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}&\text{if $m\geq 3$ is odd,}\end{cases}$ (1.4)
where $\boldsymbol{E}_{m}$ is the harmonic frame in ${\mathbb{R}}^{2}$ which is defined by
$\boldsymbol{E}_{m}:=\left(\begin{array}[]{cccc}1&\cos\frac{1}{m}\pi&\cdots&\cos\frac{m-1}{m}\pi\\ 0&\sin\frac{1}{m}\pi&\cdots&\sin\frac{m-1}{m}\pi\end{array}\right)^{T}\in\mathbb{R}^{m\times 2}.$ (1.5)
Combining (1.3) and (1.4), we have $\beta_{\boldsymbol{E}_{m}}^{\mathbb{R}}=\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}^{\mathbb{R}}$ for each odd integer $m\geq 3$. The following conjecture is proposed in [32] for each even integer $m\geq 4$.
@@
$\beta_{\boldsymbol{E}_{m}}^{\mathbb{R}}=\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}^{\mathbb{R}}.$
In this paper, we will show that minimizing the condition number of $\boldsymbol{A}\in\mathbb{R}^{m\times 2}$ can be transformed into solving the perimeter-maximizing isodiametric problem proposed by Reinhardt in 1922 [25]. We prove that each optimal solution to the perimeter-maximizing isodiametric problem gives rise to a matrix with the minimal condition number. Based on that, we can show that Conjecture 1.1 is false. For instance, when $m=4$, we will demonstrate in Remark 1.3 that
$\beta_{\boldsymbol{E}_{4}}^{\mathbb{R}}={\sqrt{2+\sqrt{2}}}\approx 1.84>\beta_{\boldsymbol{E}_{4}^{\prime}}^{\mathbb{R}}=\min_{\boldsymbol{M}\in\mathbb{R}^{4\times 2}}\beta_{\boldsymbol{M}}^{\mathbb{R}}=\sqrt{1+\frac{\sqrt{2}}{2}+\frac{\sqrt{6}}{2}}\approx 1.71,$
@@
$\boldsymbol{E}_{4}^{\prime}=\left(\begin{array}[]{cccc}-1&\cos\frac{2\pi}{3}&\sqrt{2\sin\frac{\pi}{12}}\cos\frac{3\pi}{8}&\sqrt{2\sin\frac{\pi}{12}}\cos\frac{7\pi}{24}\\ 0&\sin\frac{2\pi}{3}&\sqrt{2\sin\frac{\pi}{12}}\sin\frac{3\pi}{8}&\sqrt{2\sin\frac{\pi}{12}}\sin\frac{7\pi}{24}\end{array}\right)^{T}.$ (1.6)
@@ 1.2. The perimeter-maximizing isodiametric problem for convex polygons
The perimeter-maximizing isodiametric problem has a long and rich history. It was first studied by Reinhardt in 1922 [25], and has been furthered investigated by Vincze [30] in 1950 and Datta [14] in 1997 (see also [15, 16, 24, 9]). The convex polygon in $\mathbb{R}^{2}$ with $m$ edges is called the convex $m$-gon. The diameter of a polygon $P$, denoted by $\operatorname{diam}({P})$, is the largest distance between any pair of its vertices, i.e., $\operatorname{diam}({P}):=\max_{\boldsymbol{x},\boldsymbol{y}\,\in{P}}{\|\boldsymbol{x}-\boldsymbol{y}\|_{2}}$. The perimeter of a polygon $P$, denoted by $\operatorname{perim}({P})$, is the sum of the length of its edges. The perimeter-maximizing isodiametric problem asks:
@@
A finite set in $\mathbb{R}^{d}$ is called a frame if the vectors in this set can span $\mathbb{R}^{d}$. We say that a frame $\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathbb{R}^{d}$ is optimal if its associated matrix $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}\in\mathbb{R}^{m\times d}$ has the minimal condition number, i.e., $\beta_{\boldsymbol{A}}^{\mathbb{R}}=\min_{\boldsymbol{M}\in\mathbb{R}^{m\times d}}\beta_{\boldsymbol{M}}^{\mathbb{R}}$. A frame $\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathbb{R}^{d}$ is called a tight frame if $\sum_{i=1}^{m}\boldsymbol{a}_{i}\boldsymbol{a}_{i}^{T}=C\cdot\boldsymbol{I}_{d}$ for some constant $C>0$.
@@
$f(\boldsymbol{a})\,\,=\,\,t^{2}\cdot(\cos 2\phi,\sin 2\phi)^{T}$ (1.10)
for any $\boldsymbol{a}=t\cdot(\cos\phi,\sin\phi)^{T}\in\mathcal{T}$.
@@
We will now show that the map $f:\mathcal{T}\to\mathbb{R}^{2}$ is a bijection. The bijectivity can be established by considering two cases. Firstly, for the zero vector, it holds that $\boldsymbol{a}=\boldsymbol{0}$ if and only if $f(\boldsymbol{a})=\boldsymbol{0}$. Secondly, for any nonzero vector $\boldsymbol{e}_{0}\in\mathbb{R}^{2}$, it possesses a unique representation in the form $\boldsymbol{e}_{0}=t_{0}\cdot(\cos{2\phi_{0}},\sin{2\phi_{0}})^{T}$, where $t_{0}>0$ and $\phi_{0}\in(0,\pi]$. Given this, the unique pre-image $\boldsymbol{a}_{0}\in\mathcal{T}$ such that $f(\boldsymbol{a}_{0})=\boldsymbol{e}_{0}$ is given by $\boldsymbol{a}_{0}=\sqrt{t_{0}}\cdot(\cos{\phi_{0}},\sin{\phi_{0}})^{T}$. Therefore, $f$ is both injective and surjective, implying it is a bijection. This ensures that the inverse map $f^{-1}:\mathbb{R}^{2}\to\mathcal{T}$ is well-defined, and is explicitly given by $f^{-1}(\boldsymbol{e}_{0})=\boldsymbol{a}_{0}$.